OpenStudy (amtran_bus):

How would you rearrange this formula? Q=mCT to get the m where the q is? And the q on the other side

OpenStudy (anonymous):

First, divide both sides by CT: Q/CT = mCT/CT Q/CT = m Then just use the reflexive property to show this as m = Q/CT